OpenStudy (anonymous):

Graph each by using x and y intercepts 1. 5-y=-3 2. x+y=4 3. Y=8-6x

OpenStudy (anonymous):

To do #1, you just need to isolate y to find the y-int. So -y=-3-5, -y=-8, y=8 2: Probably change this to y=mx+b, so y=-x+4, where 4 is your y-int 3. In #3, Y-int is 8 Then, to find x-ints, you just need to set it to 0 in all of those equations.

OpenStudy (anonymous):

Thank you